OpenStudy (anonymous):

calculate the probability of randomly guessing at least 7 correct answers on a 10 questions true or false quiz to get a passing grade?

OpenStudy (anonymous):

Remember the probability of k successes in a binomial trial \[ g(k)=P[X=k]= \frac {n!}{k! \, (n-k)!} p^k (1-p)^{n-k} \] Here \[ p=q=1/2; \]

OpenStudy (anonymous):

So in this case \[ g(k) =\frac 1 {2^{10}} \frac {10!}{k! (10 -k)!} \] The answer should be \[ g(7)+g(8)+g(9) + g(10) \]

OpenStudy (anonymous):

\[ \frac{15}{128}+ \frac{45}{1024}+ \frac{5}{512}+ \frac{1}{1024} =\frac{11}{64} \]